OpenStudy (anonymous):

5/6-2/5 in simplest form. Please explain it.

OpenStudy (anonymous):

13/30

OpenStudy (anonymous):

5/6-2/5 =25/30-12/30

OpenStudy (anonymous):

=13/30

OpenStudy (anonymous):

you just take 5/6 and 2/5 and find a common denominator which in this case is 30. So 5*6 and what ever you do to the bottom you do to the top. So you get 25/30-12/30 and you get 13/30. Got it?